# Geometric Framework for Cell Oversegmentation

### v2

@@ Abstract
3D cell segmentation methods are often hindered by oversegmentation, where a single cell is incorrectly split into multiple fragments. This degrades the final segmentation quality and is notoriously difficult to resolve, as oversegmentation errors often resemble natural gaps between adjacent cells. Our work makes two key contributions. First, for 3D cell segmentation, we are the first work to formulate oversegmentation as a concrete problem and propose a geometric framework to identify and correct these errors. Our approach builds a pre-trained classifier using both 2D geometric and 3D topological features extracted from flawed 3D segmentation results. Second, we introduce a novel metric, Geo-Wasserstein divergence, to quantify changes in 2D geometries. This captures the evolving trends in cell mask shape changes in a geometry-aware manner. We validate our method through extensive experiments on in-domain plant datasets, including both synthesized and real cases, as well as on out-of-domain animal datasets to demonstrate transfer learning performance. An ablation study further highlights the contribution of the Geo-Wasserstein divergence. A clear pipeline is provided for end-users to build pre-trained models to any labeled dataset.
@@ 1 Introduction
Current 3D cell segmentation approaches can be broadly categorized into 2D-based and non-2D-based methods. Oversegmentation errors could appear differently depending on which category the segmentation approach falls under. The 2D-based methods reconstruct the 3D cell structure by “stitching” pre-segmented 2D slices layer-by-layer. This strategy has gained popularity recently due to its strong generalizability across diverse cell shapes and datasets, effectively eliminating the need for extensive pre-training on labeled 3D datasets. However, 2D-based approaches heavily rely on the quality of initial 2D segmentation results; thus, errors in individual slices can propagate and accumulate, resulting in oversegmentations predominantly along the imaging axis. In contrast, non-2D-based methods directly reconstruct 3D cell structures from raw microscopy images without using 2D pre-segmentated results. While this strategy mitigates some errors inherent in 2D segmentation, it requires substantial pre-training on large, dataset-specific annotations, thereby limiting its generalizability. Consequently, oversegmentation errors in non-2D-based methods typically manifest as arbitrary, irregular partitions cutting through the reconstructed cell bodies at random orientations.
To the best of our knowledge, there is currently no existing literature that explicitly addresses the nature of the oversegmentation problem. This challenge is particularly difficult because such oversegmentation errors closely resemble natural gaps separating adjacent cells.
In this work, we address the oversegmentation challenge by reformulating it as a binary classification problem, where the learning task is to distinguish oversegmentation from natural gaps between cells. We propose a geometric framework that quantitatively captures layer-to-layer 2D geometric consistency as well as overall 3D topological consistency across segmented cell bodies to reliably detect oversegmentation. In addition, we build a pre-trained classifier with the extracted geometric and topological information from flawed 3D segmentation results and then correct these identified oversegmented cell bodies through a proposed stitching method. We conduct extensive experiments across multiple plant and animal datasets to illustrate the effectiveness of our method. Our main contributions are summarized below:
(i) We formally study the oversegmentation issue in 3D cell segmentation as a specific learning problem and introduce a geometric framework for correcting oversegmentation errors. Our framework improves overall 3D segmentation quality and is compatible with various existing segmentation methods.
(ii) We introduce a novel geometry-aware metric, Geo-Wasserstein divergence, that leverages earthmover’s distance (EMD) to quantify the divergence between 2D geometric shapes of cell masks.
@@ 2 Preliminaries and Related Works
2.1 3D Cell Segmentation
@@
The development of 3D cell segmentation methods has been greatly facilitated by deep learning architectures such as U-Net [20], which have given rise to several state-of-the-art 2D segmentation methods, including CellPose2D [17, 25, 26], Mesmer [14], and StarDist [23]. Building on these advances, 3D segmentation methods can be broadly categorized into two types: (i) 2D-based methods, which typically rely on unsupervised or weakly supervised 2D segmentation followed by subsequent 3D reconstruction, thereby benefiting from reduced complexity and strong generalizability; and (ii) Non-2D-based methods, which directly reconstruct 3D structures from raw image stacks, generally requiring fully supervised training on large labeled datasets. Within it, 2.5D-based approaches have emerged, which integrates the 2D pre-segmented information into the final 3D contextual analysis. Such hybrid framework avoids the need for large-scale 3D pretraining while leveraging the strong generalizability of 2D-based methods
CellPose [26] (2.5D-based), CellStitch [15] (2D-based), and PlantSeg [33] (non-2D-based) represent the state-of-the-art approaches in their respective categories. However, 3D-based or deep learning-based methods [8, 28, 31] such as PlantSeg often suffers from limited generalizability due to its reliance on supervised pretraining with large labeled datasets. We review the method above and provide the review for alternative approaches in Appendix A:
@@
CellPose. CellPose is 2.5D-based approach that integrates the strengths of both 2D and 3D approaches by leveraging contextual layer awareness to develop a transferable model. It trains models to predict flow vectors for each pixel, and to generate 3D flow vectors, it averages the 2D flow vectors along the , , and planes. However, the substantial difference in sampling ratios between the plane and the -axis introduces noise, leading to oversegmentation in highly anisotropic images.
@@
CellStitch. CellStitch is an unsupervised 2D-based approach that reconstructs 3D segmentation from CellPose2D’s pre-segmentated 2D results. It uses optimal transport to match cells between adjacent layers and introduces an interpolation method to “stitch” the discrete, layer-by-layer 2D cell masks into a continuous 3D cell body. However, the performance of such 2D-based methods heavily depends on the quality of the initial 2D segmentation. Specifically, CellPose2D would mis-segment with empty or undersegmented mask in 2D layers and causing oversegmentations in CellStitch’s 3D results.
@@
PlantSeg. PlantSeg is a supervised non-2D-based method that employs deep learning to predict cell boundaries from 2D image stacks, followed by 3D reconstruction of cell bodies. Its segmentation quality is highly data-driven, depending on the image quality and the similarity between the input images and the training data. Specifically, PlantSeg is trained on labeled Arabidopsis thaliana ovule cells, and we empirically found that oversegmentation typically occurs in elongated cells.
@@ 2.2 Optimal Transport
Optimal transport (OT) has been applied to mutiple biological areas, including single-cell studies [22], drug perturbations [5, 9, 10], and cell alignments [11]. In our framework, OT is utilized to quantify the geometric change between 2D cell masks from adjacent layers. We introduce OT formulation along with earthmover’s distance (EMD) below.
@@
Eq (2.2), known as the Kantorovich formulation of optimal transport, generalizes the Monge formulation [16] by relaxing the requirement of a direct one-to-one mapping between source and target elements. For two discrete distributions and , the earthmover’s distance is defined as the minimum cost required to transform into :
@@ 3.1 Problem Formulation
We first address oversegmentation along the standard axis in 2D- and 2.5D-based models, and then generalize our approach to handle tilted cut planes at arbitrary angles. As illustrated in Figure 3, when a missing cell slice exists in layer (leaving an empty cell mask at layer ), 2D-based models would incorrectly split the entire cell into two different cells in upper section and lower section. Our task is to: (i) identify a list of suspected oversegmented candidates, consisting of upper and lower parts of cells separated by a gap; (ii) determine whether each candidate represents true oversegmentation or a natural gap, as commonly observed in loosely packed tissues like leaves, where intercellular spaces are frequent; and (iii) recover the correct 2D segmentation by predicting the cell mask for the gap layer and reconstructing the accurate 3D segmentation result. We provide a detailed explanation of these steps in the following section.
@@ 3.2 Method Pipeline
Candidates Screening. We begin by identifying suspected oversegmented candidates within the flawed 3D segmentation result. As shown in Algorithm 1, for each cell, we store its mask at the “top” (highest layer) and “bottom” (lowest layer). We then iterate over all cell pairs in the dataset. If the top-layer mask of one cell overlaps with the bottom-layer mask of another—i.e., they occupy the same position across adjacent layers—and no other cell exists between them, the pair is marked as a suspected oversegmentation candidate.
Binary Classification. For each suspected candidate, the goal is to distinguish true oversegmentations from natural gaps. We frame this as a binary classification task and train a classifier using features extracted from labeled examples of both oversegmented and natural-gap cell pairs.
@@
Geometric Features Extraction. For each candidate pair, we quantify the geometric change at the gap layer and compare it to the typical layer-to-layer transitions within each cell. In true oversegmentations, the gap-layer change tends to be smooth and consistent with prior transitions. In contrast, an abrupt or irregular change at the gap suggests the presence of two distinct cells, indicating a natural separation. For cell masks and , we represent them as uniform distributions over their geometric regions:
@@
where is the indicator function denoting whether lies within the boundary of , and denotes the area of mask . The Geo-Wasserstein divergence (GW divergence) between these two masks—i.e., the change in geometric shape—can then be quantified using the earthmover’s distance: . To assess whether the GW divergence between two gap-layer masks is consistent with the prior trend of mask transitions in each cell, we use Algorithm 2 to extract EMD values; each cell is indexed from layer (top) to layer (bottom). Moreover, since different cells have varying heights, the lengths of the arrays differ. To enable training, it is necessary to standardize the input length across all samples. To achieve this, we extract key statistical features from each EMD array to serve as inputs to the model. Specifically, in our pretrained model, we use summary statistics such as the median, max, min, first quartile (), and third quartile () for each array. These features, together with the gap-layer EMD value , are used as inputs to the training process.
@@
Topological Shape Extraction. We propose a simple yet effective method from 3D topological perspective to distinguish oversegmented cell pairs from those separated by natural gaps. As illustrated in Figure 4, a key pattern observed across multiple datasets is that oversegmented cases typically exhibit a smooth shape transition at the gap layer, often following either a linear or circular progression. These patterns have biological relevance: circular shapes—wider in the middle and narrower at the ends—are characteristic of mitotic cells undergoing “circular rounding” [21, 6], while gradually tapering linear shapes are typical of structural cells such as palisade cells [18].
We quantify such linearity and circularity by analyzing the overlapping area between masks at adjacent layers. For cells exhibiting a linear pattern, the overlap area should change in a monotonically linear trend. In contrast, for circular patterns, the overlap area is expected to follow a strictly quadratic trend—characterized by a monotonous increase followed by a decrease. As shown in the Table 1, we evaluate this strict pattern across cells in the labeled datasets.
| Dataset | # of Cells | % SM |
| Filament | 5,879 | 0.963 |
| Leaf | 2,207 | 0.927 |
| Anther | 7,711 | 0.920 |
| Sepal | 6,744 | 0.908 |
| Pedicel | 7,319 | 0.869 |
| Valve | 15,817 | 0.851 |
@@
However, by examining the cells that deviate from the strictly monotonous pattern, we observed occasional oscillations in the overlapping area changes—despite the overall shape following a clear linear or quadratic trend (see Figure 9). To improve the inclusiveness towards more candidates, we relaxed the strict constraint by fitting linear and quadratic regressions to the overlap curve and using the regression coefficient as a measure of shape consistency. For strictly monotonous cells, we assign . Each suspected candidate pair is treated as a single cell, and we apply Algorithm 3 to extract its shape information.
Classifier Training. Therefore, by leveraging both sources of information, we aim to evaluate whether the gap layer maintains geometric consistency with the preceding layers, and whether the entire cell exhibits topological consistency typical of normal cells. Our full pipeline for classifier training is illustrated in Figure 2.
@@
We begin by constructing a training set composed of labeled true (oversegmented) and false (natural gap) cases from annotated datasets. True cases are generated by simulating oversegmentation—removing a specific layer from a 2D segmentation to create an artificial gap. False cases are obtained by applying Algorithm 1 to identify naturally separated cells. We then apply Algorithm 2 and Algorithm 3 to each candidate in the training set to extract geometric and topological shape features. To preserve all extracted information, we concatenate the features and use them as input to a multi-layer perceptron, which is trained to perform binary classification.
@@
Cross-layer Interpolation. For final oversegmented cell pairs, we recover the mis-segmented 2D mask between them (at layer ) by interpolating between the cell mask boundaries at layers and , using the interpolation method from CellStitch [15]. The corrected 2D segmentation masks are then used to reconstruct the updated 3D segmentation. The specific implementation details of this cross-layer interpolation are provided in Appendix B.1.

### v3

# 3D Cell Oversegmentation Correction via Geo-Wasserstein Divergence

@@ Abstract
3D cell segmentation methods are often hindered by oversegmentation, where a single cell is incorrectly split into multiple fragments. This degrades the final segmentation quality and is notoriously difficult to resolve, as oversegmentation errors often resemble natural gaps between adjacent cells. Our work makes two key contributions. First, for 3D cell segmentation, we are the first work to formulate oversegmentation as a concrete problem and propose a geometric framework to identify and correct these errors. Our approach builds a pre-trained classifier using both 2D geometric and 3D topological features extracted from flawed 3D segmentation results. Second, we introduce a novel metric, Geo-Wasserstein divergence, to quantify changes in 2D geometries. This captures the evolving trends of cell mask shape in a geometry-aware manner. We validate our method through extensive experiments on in-domain plant datasets, including both synthesized and real oversegmented cases, as well as on out-of-domain animal datasets to demonstrate transfer learning performance. An ablation study further highlights the contribution of the Geo-Wasserstein divergence. A clear pipeline is provided for end-users to build pre-trained models to any labeled dataset.
@@ 1 Introduction
Current 3D cell segmentation approaches can be broadly categorized into 2D-based and non-2D-based methods. Oversegmentation errors could appear differently depending on which category the segmentation approach falls under. The 2D-based methods reconstruct the 3D cell structure by “stitching” pre-segmented 2D slices layer-by-layer. This strategy has gained popularity recently due to its strong generalizability across diverse cell shapes and datasets, effectively eliminating the need for extensive pre-training on labeled 3D datasets. However, 2D-based approaches heavily rely on the quality of initial 2D segmentation results; thus, errors in individual slices can propagate and accumulate, resulting in oversegmentations predominantly along the imaging axis. In contrast, non-2D-based methods directly reconstruct 3D cell structures from raw image stacks without using 2D pre-segmented results. While this strategy mitigates some errors inherent in 2D segmentation, it requires substantial pre-training on large, dataset-specific annotations, thereby limiting its generalizability. Consequently, oversegmentation errors in non-2D-based methods typically manifest as arbitrary, irregular partitions cutting through the reconstructed cell bodies at random orientations.
To the best of our knowledge, there is currently no existing literature that explicitly addresses the nature of the oversegmentation problem. This challenge is particularly difficult because oversegmentation errors closely resemble natural gaps separating adjacent cells. In this work, we address the oversegmentation challenge by reformulating it as a binary classification problem, where the learning task is to distinguish oversegmentation from natural gaps between cells. We propose a novel geometric framework that quantitatively captures layer-to-layer 2D geometric consistency as well as overall 3D topological consistency across segmented cell bodies to reliably detect oversegmentation. In addition, we build a pre-trained classifier with the extracted geometric and topological information from flawed 3D segmentation results and then correct these identified oversegmented cell bodies through a proposed interpolation method. Our main contributions are summarized below:
(i) We formally study the oversegmentation issue in 3D cell segmentation as a specific learning problem and introduce a simple yet effective framework for correcting oversegmentation errors. Our framework improves overall 3D segmentation quality and is compatible with various existing state-of-the-art segmentation methods.
(ii) We introduce a novel geometry-aware metric, Geo-Wasserstein divergence, that leverages earthmover’s distance to quantify the divergence between 2D geometric shapes of cell masks. We further highlight its potential application to broader scenarios in general segmentation task for the similar trait of temporal consistency.
@@ 2 Preliminaries and Related Works
2.1 3D Cellular Anisotropic Image Segmentation
@@
The development of cell segmentation has been greatly facilitated by deep learning architectures such as U-Net [22], which have given rise to several state-of-the-art 2D segmentation methods, including CellPose2D [19, 27, 28], Mesmer [16], and StarDist [25]. Building on these advances, 3D segmentation methods can be broadly categorized into two types: (i) 2D-based methods, which typically rely on unsupervised or weakly supervised 2D segmentation followed by subsequent 3D reconstruction, thereby benefiting from reduced complexity and strong generalizability; and (ii) Non-2D-based methods, which directly reconstruct 3D structures from raw image stacks, generally requiring fully supervised training on large labeled datasets. Within it, 2.5D-based approaches have emerged, which integrates the 2D pre-segmented information into the final 3D contextual analysis. Such hybrid framework avoids the need for large-scale 3D pretraining while leveraging the strong generalizability of 2D-based methods. CellPose [28] (2.5D-based), CellStitch [17] (2D-based), and PlantSeg [35] (non-2D-based) represent the state-of-the-art approaches in their respective categories. However, 3D-based or deep learning-based methods [8, 31, 32] such as PlantSeg often suffers from limited generalizability due to its reliance on supervised pretraining with large labeled datasets. We review the method above and provide a technical review for more 2D segmentation backbones [22, 15, 21, 7, 13, 36] and 3D segmentation approaches [30, 1, 33, 29] in Appendix A.
CellPose.
@@
CellPose is a 2.5D-based approach that integrates the strengths of both 2D and 3D approaches by leveraging contextual layer awareness to develop a transferable model. It trains models to predict flow vectors for each pixel, and to generate 3D flow vectors, it averages the 2D flow vectors along the , , and planes. However, the substantial difference in sampling ratios between the plane and the -axis introduces noise, leading to oversegmentation in highly anisotropic images.
CellStitch.
@@
CellStitch is an unsupervised 2D-based approach that reconstructs 3D segmentation from CellPose2D’s pre-segmentated 2D results. It uses optimal transport to match cells between adjacent layers and introduces an interpolation method to “stitch” the discrete, layer-by-layer 2D cell masks into a continuous 3D cell body. However, the performance of such 2D-based methods heavily depends on the quality of the initial 2D segmentation. Specifically, CellPose2D would mis-segment with empty or undersegmented mask in 2D layers and causing oversegmentations in CellStitch’s 3D results.
PlantSeg.
@@
PlantSeg is a supervised non-2D-based method that employs deep learning to predict cell boundaries from 2D image stacks, followed by 3D reconstruction of cell bodies. Its segmentation quality is highly data-driven, depending on the image quality and the similarity between the input images and the training data.
@@ 2.2 Optimal Transport
Optimal transport (OT) has been applied to mutiple biological areas, including single-cell studies [24], drug perturbations [5, 10, 11, 9], and cell alignments [12]. In our framework, OT is utilized to quantify the geometric change between 2D cell masks from adjacent layers. We introduce OT formulation along with earthmover’s distance (EMD) below.
@@
Formulation (2.2), known as the Kantorovich formulation of optimal transport, generalizes the Monge formulation [18] by relaxing the requirement of a direct one-to-one mapping between source and target elements. For two discrete distributions and , the earthmover’s distance is defined as the minimum cost required to transform into :
@@ 3.1 Problem Formulation
We first address oversegmentation along the standard axis in 2D- and 2.5D-based models, and then generalize our approach to handle tilted cut planes at arbitrary angles. As illustrated in Figure 3, when a missing cell slice exists in layer (leaving an empty cell mask at layer ), 2D-based models would incorrectly split the entire cell into two different cells in upper section and lower section. Our task is to: (i) identify a list of suspected oversegmented candidates, consisting of upper and lower parts of cells separated by a gap; (ii) determine whether each candidate represents true oversegmentation or a natural gap, as commonly observed in loosely packed tissues like leaves, where intercellular spaces are frequent; and (iii) recover the correct 2D segmentation by predicting the cell mask for the gap layer and reconstructing the accurate 3D segmentation result. We provide a detailed pipeline in following sections.
@@ 3.2 Method Pipeline
We begin by identifying suspected oversegmented candidates within the flawed 3D segmentation result. As shown in Algorithm 1, for each cell, we store its mask at the “top” (highest layer) and “bottom” (lowest layer). We then iterate over all cell pairs in the dataset. If the top-layer mask of one cell overlaps with the bottom-layer mask of another—i.e., they occupy the same position across adjacent layers—and no other cell exists between them, the pair is marked as a suspected oversegmentation candidate. For the suspected candidates collected from candidate screening stage, our goal is to differentiate real oversegmented cases from the case with natural gap between two adjacent cells. Before delving into the technical details of feature extraction, we first introduce some biological background on the cells used in 3D reconstruction to provide better context for our method.
Cells used in 3D Segmentation.
3D segmentation is routinely applied to extract quantitative morphometric information (volumes, neighbour counts, and growth anisotropy) from an entire organ or tissue that has been imaged as a stack. This workflow requires that every voxel from each layer be assigned unambiguously to a single cell, which in practice excludes cells whose highly irregular shapes create multiple, overlapping masks within the same plane. Moreover, standard confocal stacks are anisotropic (i.e., ). Meanwhile, many irregular cell types (e.g., neuron dendrite) that lie outside compact organs are thinner than the ‑sampling interval and appear discontinuous in the volume. Hence, such complex morphologies are usually analyzed with connectomics pipelines based on electron microscopy, rather than with the volumetric segmentation methods for our work. We therefore consider the cells from organ-level image stacks used in 3D reconstruction, assuming a certain level of regularity that can be captured through smooth geometric transitions at temporal level.
Through extensive experiments on both plant and animal cells, we validate this assumption in the context of 3D cellular image segmentation, which forms the basis for the subsequent geometric feature extraction framework using Geo-Wasserstein divergence.
Geometric Features Extraction.
@@
For each candidate pair, we quantify the geometric change at the gap layer and compare it to the typical layer-to-layer transitions within each cell. In true oversegmentations, the gap-layer change tends to be smooth and consistent with prior transitions. In contrast, an abrupt or irregular change at the gap suggests the presence of two distinct cells, indicating a natural separation. For cell masks and , we represent them as uniform distributions over their geometric regions:
@@
where is the indicator function denoting whether lies within the boundary of , and denotes the area of mask . The Geo-Wasserstein divergence between these two masks—i.e., the change in geometric shape—can then be quantified using the earthmover’s distance: . To assess whether the Geo-Wasserstein divergence between two gap-layer masks is consistent with the prior trend of mask transitions in each cell, we use Algorithm 2 to extract EMD values; each cell is indexed from layer (top) to layer (bottom).
Moreover, since different cells have varying heights, the lengths of the arrays differ. To enable training, it is necessary to standardize the input length across all samples. To achieve this, we extract key statistical features from each EMD array to serve as inputs to the model. Specifically, in our pretrained model, we use summary statistics such as the median, max, min, first quartile (), and third quartile () for each array. These statistics are designed to capture higher-level trends in cell shape-change trajectories, which serves as an effective strategy to account for the varying cell heights. We further conduct ablation analysis in §4.4 for these statistics to justify the robustness and sensitivity. These features, together with the gap-layer EMD value , are used as inputs to the training process.
Topological Shape Extraction.
@@
We propose a straightforward method from 3D topological perspective to distinguish oversegmented cell pairs from those separated by natural gaps. As illustrated in Figure 4, a key pattern observed across multiple datasets is that oversegmented cases typically exhibit a smooth shape transition at the gap layer, often following either a linear or circular progression. These patterns have biological relevance: circular shapes—wider in the middle and narrower at the ends—are characteristic of mitotic cells undergoing “circular rounding” [23, 6], while gradually tapering linear shapes are typical of structural cells such as palisade cells [20]. We quantify such linearity and circularity by analyzing the overlapping area between masks at adjacent layers. For cells exhibiting a linear pattern, the overlap area should change in a monotonically linear trend. While for circular patterns, the overlap area is expected to follow a strictly quadratic trend, characterized by a monotonous increase followed by a decrease. As shown in the Table 1, we evaluate this strict pattern across cells in the labeled datasets:
| Dataset | Filament | Leaf | Anther | Sepal | Pedicel | Valve |
| # of Cells | 5,879 | 2,207 | 7,711 | 6,744 | 7,319 | 15,817 |
| % SM | 0.963 | 0.927 | 0.920 | 0.908 | 0.869 | 0.851 |
@@
However, by examining the cells that deviate from the strictly monotonous pattern, we observed occasional oscillations at specific layers, despite the overall shape following a clear linear or quadratic trend. To improve the inclusiveness towards more candidates, we relaxed the strict constraint by fitting linear and quadratic regressions to the overlap curve and using the normalized regression coefficient as a measure of shape consistency. For strictly monotonous cells, we assign . Each suspected candidate pair is treated as a single cell, and we apply Algorithm 3 to extract its shape information.
Classifier Training.
@@
By leveraging both sources of information, we aim to evaluate whether the gap layer maintains geometric consistency with the preceding layers, and whether the entire cell exhibits topological consistency typical of normal cells. Our full pipeline for classifier training is illustrated in Figure 2. We begin by constructing a training set composed of labeled true (oversegmented) and false (natural gap) cases from annotated datasets. True cases are generated by simulating oversegmentation—removing a specific layer from a 2D segmentation to create an artificial gap. False cases are obtained by applying Algorithm 1 to identify naturally separated cells. We then apply Algorithm 2 and Algorithm 3 to each candidate in the training set to extract geometric and topological shape features. To preserve all extracted information, we concatenate the features and use them as input to a multi-layer perceptron, which is trained to perform binary classification.
Cross-layer Interpolation.
@@
For final oversegmented cell pairs, we recover the mis-segmented 2D mask between them (at layer ) by interpolating between the cell mask boundaries at layers and , leveraging the interpolation methods [17, 26]. The corrected 2D segmentation masks are then used to reconstruct the updated 3D segmentation. The technical details of cross-layer interpolation are provided in Appendix B.1.
Computational Efficiency.
We highlight the computational efficiency of this lightweight, effective method and report timings from real applications. Processing a large animal stack takes 1 hour on two NVIDIA A40 GPUs (48GB each). On a medium plant dataset, it takes 7 minutes to identify oversegmentation candidates and 15 minutes to stitch and recover all oversegmented cells, using CPUs only.